# Goal-Conditioned Neural ODEs with Guaranteed Safety and Stability for Learning-Based All-Pairs Motion Planning

### 03 Apr 2026 v1

@@ III-B Safety Properties
The key property our approach is that it guarantees safety for arbitrary start/goal pairs. In the control literature, safety in the form of forward invariance of a set is often certified by a barrier function, as defined below (see e.g. [36]):
@@
$\begin{split}\dot{h}&=2\lambda(\lambda|g(x)|^{2}-g(x)^{\top}g(x_{\star}))\\ &\geq 2\lambda|g(x)|(|g(x)|-|g(x_{\star}|)>0.\end{split}$ (22)
@@ III-C Time-Varying Goal Location
    $|\epsilon(t)|\leq\frac{\nu}{\mu}\left(|\epsilon(0)|e^{-\lambda t}+b\right).$ (23)
@@
$\dot{\epsilon}_{z}=-\lambda\epsilon_{z}+\lambda\dot{z}_{\star}(t),$
where $|\dot{z}_{\star}(t)|\leq\nu b$. This implies $|\epsilon_{z}(t)|\leq|\epsilon_{z}(0)|e^{-\lambda t}+\nu b$. Finally, following the procedure in (18) yields
$\begin{split}|\epsilon(t)|\leq\frac{1}{\mu}|\epsilon_{z}(t)|\leq\frac{\nu}{\mu}\left(|\epsilon(0)|e^{-\lambda t}+b\right).\end{split}$
@@ III-D Finite-time Convergence via Euclidean Norm Potential
From [19, Thm. 1] we can conclude that the above system achieves equilibrium-independent exponential stability. However, it cannot provide safety guarantees for all $x_{\star}\in\mathcal{X}_{\text{safe}}$, see the example below.
@@ III-G Compared with Existing Diffeomorphism based Dynamical Approaches
where the potential function $\Phi$ is positive definite, convex, convex, continuously differentiable, and radially unbounded. And a natural gradient dynamics in the $\mathcal{X}$-space is constructed by pulling (31) back to the $\mathcal{X}_{\text{safe}}$-space via a diffeomorphism $g$. The primary goal of those approaches is to learn stable dynamics that mimics the demonstration data.

### 14 Apr 2026 v2

@@ III-B Safety Properties
The key property of our approach is that it guarantees safety for arbitrary start/goal pairs. In the control literature, safety in the form of forward invariance of a set is often certified by a barrier function, as defined below (see e.g. [36]):
@@
$\begin{split}\dot{h}&=2\lambda(|g(x)|^{2}-g(x)^{\top}g(x_{\star}))\\ &\geq 2\lambda|g(x)|(|g(x)|-|g(x_{\star})|)>0.\end{split}$ (22)
@@ III-C Time-Varying Goal Location
    $|\epsilon(t)|\leq\frac{\nu}{\mu}\left(|\epsilon(0)|e^{-\lambda t}+\frac{b}{\lambda}\right).$ (23)
@@
$\dot{\epsilon}_{z}=-\lambda\epsilon_{z}-\dot{z}_{\star}(t),$
where $|\dot{z}_{\star}(t)|\leq\nu b$. This implies $|\epsilon_{z}(t)|\leq|\epsilon_{z}(0)|e^{-\lambda t}+\nu b/\lambda$. Finally, following the procedure in (18) yields
$\begin{split}|\epsilon(t)|\leq\frac{1}{\mu}|\epsilon_{z}(t)|\leq\frac{\nu}{\mu}\left(|\epsilon(0)|e^{-\lambda t}+\frac{b}{\lambda}\right).\end{split}$
@@ III-D Finite-time Convergence via Euclidean Norm Potential
From [19, Thm. 1], we can conclude that the above system achieves equilibrium-independent exponential stability. However, it cannot provide safety guarantees for all $x_{\star}\in\mathcal{X}_{\text{safe}}$, see the example below.
@@ III-G Compared with Existing Diffeomorphism based Dynamical Approaches
where the potential function $\Phi$ is positive definite, convex, continuously differentiable, and radially unbounded. And a natural gradient dynamics in the $\mathcal{X}$-space is constructed by pulling (31) back to the $\mathcal{X}_{\text{safe}}$-space via a diffeomorphism $g$. The primary goal of those approaches is to learn stable dynamics that mimics the demonstration data.